OpenStudy (anonymous):

y=-(2/3)x+30 trying to find x and y intercept

jimthompson5910 (jim_thompson5910):

x-intercept: plug in y=0 and solve for x y=-(2/3)x+30 0=-(2/3)x+30 -30=-(2/3)x -30*3=-2x -90=-2x -90/(-2)=x 45=x x=45 So the x-intercept is (45,0) -------------------- y-intercept: plug in x=0 and solve for y y=-(2/3)x+30 y=-(2/3)(0)+30 y=0+30 y=30 So the y-intercept is (0,30)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

when solving for y=-(2/3)x+30....you did you get -30*3?

jimthompson5910 (jim_thompson5910):

not sure what you mean, are you referring to my 4th step?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

in going from -30=-(2/3)x to -30*3=-2x I multiplied both sides by 3. Why? Because on the right side we're dividing by 3. So we undo that to solve for x by multiplying both sides by 3.

OpenStudy (anonymous):

sorry could you go one step back ? how did you get the -30 in front of the = sign. I'm horrible in math!

jimthompson5910 (jim_thompson5910):

I went from 0=-(2/3)x+30 to -30=-(2/3)x by subtracting 30 from both sides

OpenStudy (anonymous):

why?

jimthompson5910 (jim_thompson5910):

because I want to solve for x

jimthompson5910 (jim_thompson5910):

ie get x on one side all by itself

OpenStudy (anonymous):

OH! thats right you want X by it self correct?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

WOW! Thanks alot just got it!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

sorry one more question? why did you multiply 3to -30 in -(2/3) why the 3?

jimthompson5910 (jim_thompson5910):

because /3 means "divide by 3". So undo that and solve for x, we multiply both sides by 3

OpenStudy (anonymous):

so what is it doesn't divide evenly?

jimthompson5910 (jim_thompson5910):

that part doesn't matter. All I want to do is move that 3 to the other side and I do so by multiplying both sides by 3

OpenStudy (anonymous):

k. thanks